Assume a town where there are only two newspaper: The globe and the mail. 60% of the people read the globe, 80% read the mail, and 10% read neither newspaper. Calculate the probability that a randomly chosen person reads (a) Both newspapers (B) the globe but not the mail (c) the mail if it is known that the chosen person does not read the globe, (d) no more than one newspaper and (e) does not read the globe.
Answer #1

Let A and B denote the events that a person reads the globe and the mail respectively.

Then,

a)

Now,

We know, which is the required probability.

b)

Required probability =

c)

Required probability =

d)

Required probability = P(no more than one newspaper) = 1 - P(both newspaper) =

e)

Required probability =